I know how to set the problem up but I do not understand how they the get the answer to this problem:
How many ways can seven books be arranged on a book shelf?
The answer is: 7!/1=5,040

Answer by stanbon(75887)

How many ways can seven books be arranged on a book shelf?
The answer is: 7!/1=5,040
----
The answer is "the permutation of 7 objects taking all 7".
-------
7P7 = 7!/(7-7)! = 7!/1 = 7*6*5*4*3*2^1= 5040
======================
Cheers,
Stan H.
=============